Firm B High Price Low Price High Price A:8 B:8 A;3 B:20 Low Price A:20 B:3...

Firm B

High Price Low Price

High Price

A:8

B:8

A;3

B:20

Low Price

A:20

B:3

A:6

B:6

*Firm A on Left, Firm B on top

- What is the PV of profits if the firms collude successfully assuming an interest rate of 15%?

- What is PV of profits for Firm A if it defects from the collusive behavior assuming an interest rate of 15%?
Answer #1

Discount factor d = 1/1 + r

d = 1/1.15 = 0.8696

Present value of profits for any of the firm when they collude (assuming infinite time period) = 8. Total profit for both firms = 16.

Present value of profits for the defecting firm = 1 period profit(1 - d) + defecting profit x d

= 20(1 - 0.8696) + 6*0.8696

= 7.82.